If I have one gallon of product wih an active ingredient of 5% and would like to dilute the product with water to make a solution with .75% active ingredient, how many gallons of water would I need to add?
Found 2 solutions by Alan3354, solver91311:
Answer by Alan3354(69443)

If I have one gallon of product wih an active ingredient of 5% and would like to dilute the product with water to make a solution with .75% active ingredient, how many gallons of water would I need to add?
----------------
There's 0.05 gallon of active ingredient in the gallon.
Find how many gallons total would make that 0.0075 of the total.
0.05/total = 0.0075
total = 0.05/0.0075 gallons
total = 500/75 = 20/3 gallons.
Add 20/3 - 1 gallons = 17/3 gallons of water
17/3 = 5 2/3 gallons.

Answer by solver91311(24713)



Let represent the amount of water in gallons you want to add. If 5% of the gallon is active ingredient, then 95% is inactive. Also, you want to end up with 0.75 solution, so 99.25% is inactive. Note that the total amount of solution in the end will be , the amount of water added, plus 1 gallon, the original amount of 5% solution.



Just solve for


John